Assuming that boys and girls are equally likely, estimate the probability of getting more than 3600 girls in 6400 births.
Answer by stanbon(75887)

Assuming that boys and girls are equally likely, estimate the probability of getting more than 3600 girls in 6400 births.
-------
mean = np = (1/2)6400 = 3200
std = sqrt(npq) = sqrt(1600) = 40
------------------------
Find the z-score of 3600
z(3600) = (3600-3200)/40 = 400/40 = 10
---
P(x > 3600) = P(z > 10) = 7.8 x 10^(-24)
==============================================
Cheers,
Stan H.